Do some research and then describe the operation of a spring-coil galvanometer.
Answer #1

to explain the operation of a spring-coil galvanometer

if the galvanometer having N coils that carry current kept in strong magnets

due to the coil rotates the rorque will be generated

that torque is \tau = k \theta

and

\tau = N i A B

by equating

N i A B = k \theta

\theta / i = N A B / k